5. (20 points) Determine the moments of inertia of the shaded area shown with respect to the r and y axes. 0.
Answer #1

Solution: A A a X B B We have, Ix = (Ix) + (1x)2 - (1x), where, (1x), = (2.0) (29 IT 4 8 (Inn), 5(10o), (IAA)2 (is) - (in), -